A survey of 1000 men ages 20 to 30 found that their heights were normally distributed, with a mean of 65 inches and a standard deviation of 2.5 inches. How many men have a height that is between 60 inches and 70 inches?


1
Expert's answer
2022-05-05T10:49:20-0400
"P(60<X<70)=P(Z<\dfrac{70-65}{2.5})"

"-P(Z\le\dfrac{60-65}{2.5})=P(Z<2)-P(Z\le-2)"

"\approx0.97725-0.02275\approx0.9545"

"0.9545(1000)=955"

955 men.